The Table shows some old measures of liquid volume. To complete the table, what numbers should be entered in (a) the wey column, (b) the chaldron column, (c) the bag column, (d) the pottle column, and (e) the gill column? (f) The volume of 1 bag is equal to 0.1091 m3. If an old story has a witch cooking up some vile liquid in a cauldron of volume 1.7 chaldrons, what is the volume in cubic meters?
Answer #1

Part A.

From the 1st row we can see that

1 wey = 120240 gill, So

1 gill = (1/120240) wey

1 gill = 8.32*10-6 wey

Part B.

From the 2nd row, we know that

1 chaldron = 576 pottle

1 pottle = (1/576) chaldron

1 pottle = 1.74*10-3 chaldron

Part C.

From the 3rd row, we know that

1 bag = 8.33*10^-2 chaldron

1 chaldron = 1/(8.33*10^-2) bag

1 chaldron = 12.00 bag

Part D.

From the 4th row, we know that

1 pottle = 2.08*10^-2 bag

1 bag = 1/(2.08*10^-2) pottle

1 bag = 48.1 pottle

Part E.

From the 5th row, we know that

1 gill = 5.32*10^-3 pottle

1 pottle = 1/(5.32*10^-3) gill

1 pottle = 188.0 gill

Part F.

Given that

1 bag = 0.1091 m^3

1.2 chaldron = ? m^3

Now we know that

1 chaldron = 12.00 bag, (Part C) So

1.2 chaldron = 12.00*1.2 = 14.4 bag = 14.4*0.1091 m^3

1.2 chaldron = 1.57 m^3

please remove dislike you have not posted table also